# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 5.2. Expressibility of Valued Relations
It follows that $(<)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. Observe that $(<)_{0}^{1}$ equals $R_{1,0,1}$. Therefore, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ by Lemma 5.14, as we wanted to prove. If $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\min^{*}$, $\operatorname{mx}^{*}$, $\operatorname{mi}^{*}$, or $\operatorname{ll}^{*}$, we use the dual versions of Proposition 5.11 and Theorem 5.13 to analogously prove that $(>)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. Since $(<)_{0}^{1}(x,y)=(>)_{0}^{1}(y,x)$ we obtain $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ by Lemma 5.14. ∎
Lemma 5.22.
Let ${\mathfrak{A}}$ be a valued structure with $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$ and $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$. Suppose that $\operatorname{lex}\notin\operatorname{fPol}({\mathfrak{A}})$ and $\operatorname{lex}\in\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Then $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$.
Proof.
Let $R$ be a valued relation of ${\mathfrak{A}}$ of arity $k$ that is not improved by $\operatorname{lex}$. Then there exist $s,t\in\mathbb{Q}^{k}$ such that
$R(s)+R(t)<2R(\operatorname{lex}(s,t)).$
In particular, $R(s),R(t)<\infty$. Since $\operatorname{Feas}(R)\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ is improved by $\operatorname{lex}$, we have $R(\operatorname{lex}(s,t))<\infty$. Let $u:=\operatorname{lex}(s,t)$. Note that we must have $R(s)<R(u)$ or $R(t)<R(u)$. Moreover, $E_{u}=E_{s}\cap E_{t}$. Let $v\in\{s,t\}$ be such that $R(v)<R(u)<\infty$. Note that we have $E_{u}\subseteq E_{v}$. Let $O$ be a maximal subset of $O_{u}$ such that there exists $w\in\mathbb{Q}^{k}$ satisfying
@@
Since $R(w)\neq R(u)$ and $E_{u}\subseteq E_{w}$, there exist $i,j\in[k]$ such that $w_{i}\leq w_{j}$ and $u_{i}>u_{j}$. Without loss of generality me may assume $(i,j)=(1,2)$, because otherwise we permute the entries of $R$. Let
$\phi(x_{1},x_{2}):=\min_{x_{3},\dots,x_{k}}\left(R(x_{1},\dots,x_{k})+\sum_{(p,q)\in E_{u}}(=)_{0}^{\infty}(x_{p},x_{q})+\sum_{(p,q)\in O}(<)_{0}^{\infty}(x_{p},x_{q})\right).$ (7)
Let $a,b\in\mathbb{Q}$ such that $a<b$. Then $\phi(b,a)=\phi(u_{1},u_{2})\leq R(u)$, because $O\subseteq O_{u}$. Suppose that $\phi(b,a)\leq R(w)$. Then there exists $w^{\prime}\in\mathbb{Q}^{k}$ that realizes the minimum in (7) and hence $\phi(b,a)=R(w^{\prime})\leq R(w)\leq R(v)$ and $w_{1}^{\prime}>w_{2}^{\prime}$. In particular, the sums in (7) are finite. Therefore, $O\cup\{(2,1)\}\subseteq O_{w^{\prime}}$ and $E_{u}\subseteq E_{w^{\prime}}$. Since $(2,1)\in O_{u}\setminus O$, this contradicts the choice of $O$ and $w$. Therefore, $\phi(b,a)>R(w)$. Note that $\phi(w_{1},w_{2})\leq R(w)$. If $w_{1}<w_{2}$, then $\phi(a,b)\leq R(w)$ and $\phi$ expresses $R_{\alpha,\beta,\gamma}$ where $\beta=\phi(a,b)$ and $\gamma=\phi(b,a)$. In particular, $\beta<\gamma<\infty$. Therefore, by Lemma 5.15, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$. Otherwise we have $w_{1}=w_{2}$. Then $\phi(a,a)\leq R(w)$ and $\phi$ expresses $R_{\alpha,\beta,\gamma}$ where $\alpha=\phi(a,a)\leq R(w)<\phi(b,a)=\gamma$. If $\beta\geq\gamma$, then $\alpha<\min(\beta,\gamma)$, and otherwise $\beta<\gamma<\infty$. In both cases, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ by Lemma 5.15. ∎
5.3. Classification
We are now ready to prove the complexity dichotomy for temporal VCSPs. We first phrase the classification with 4 cases, where we distinguish between the tractable cases that are based on different algorithms. As a next step, we formulate two corollaries each of which provides two concise mutually disjoint conditions that correspond to NP-completeness and polynomial-time tractability, respectively.
Theorem 5.23.
@@ 5.3. Classification
    $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$ (see Definition 5.1), $-T_{3}$, or $\operatorname{Dis}$ (see (5)). In this case, ${\mathfrak{A}}$ has a reduct ${\mathfrak{A}}^{\prime}$ over a finite signature such that $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
  2. (2)
@@
    $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$. In this case, for every reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ over a finite signature, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P.
  3. (3)
    $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})$ and $\operatorname{Pol}(\hat{{\mathfrak{A}}})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. In this case, for every reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ over a finite signature, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P.
  4. (4)
    $\pi_{1}^{2}\in\operatorname{fPol}({\mathfrak{A}})$ and $\operatorname{fPol}({\mathfrak{A}})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. In this case, for every reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ over a finite signature, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P.
Proof.
@@
Note that for every reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$, the automorphism group $\operatorname{Aut}({\mathfrak{A}}^{\prime})$ contains $\operatorname{Aut}({\mathfrak{A}})$ and hence is oligomorphic. If $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$, then there is a reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ over a finite signature such that $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is NP-hard by Lemma 2.7 and Theorem 5.8. By Theorem 2.25, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in NP, therefore it is NP-complete. If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}}^{\prime})$ for every reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ over a finite signature, and $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P by Lemma 3.1. Suppose therefore that $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$ and that $\langle{\mathfrak{A}}\rangle$ does not contain any of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$. By Lemma 4.2, $\langle{\mathfrak{A}}\rangle$ contains $(\neq)_{0}^{\infty}$ or $(<)_{0}^{\infty}$, and hence $\operatorname{const}\notin\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Recall that $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ contains all relations primitively positively definable in $(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ (Remark 2.9). By Theorem 5.8, $\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ (and thus $\operatorname{Pol}(\hat{{\mathfrak{A}}})$) contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals.
@@
Let ${\mathfrak{A}}^{\prime}$ be a reduct of ${\mathfrak{A}}$ with a finite signature. If $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Sym}(\mathbb{Q})$, then by Theorem 4.6, $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})\subseteq\operatorname{fPol}({\mathfrak{A}}^{\prime})$ and $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P. By Lemma 5.6, and since $\operatorname{Aut}({\mathfrak{A}})$ contains $x\mapsto-x$, we have $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})$ and therefore satisfy (3).
@@
Finally, suppose that $\operatorname{Aut}({\mathfrak{A}})\neq\operatorname{Sym}(\mathbb{Q})$. By Lemma 5.16 we have $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$, and hence $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$. By Lemma 5.21 we have that ${\mathfrak{A}}$ is essentially crisp or $\operatorname{lex}\in\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. If ${\mathfrak{A}}$ is not essentially crisp, we have $\operatorname{lex}\in\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$, and $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})\subseteq\operatorname{fPol}({\mathfrak{A}}^{\prime})$ by Lemma 5.22. Then $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P by Lemma 5.19 and Condition (3) holds. Suppose that ${\mathfrak{A}}$ is essentially crisp. Then by Lemma 3.3 we have $\pi_{1}^{2}\in\operatorname{fPol}({\mathfrak{A}})$. Since $\operatorname{const}\not\in\operatorname{fPol}({\mathfrak{A}})$, we have $\operatorname{const}\not\in\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$ (see Remark 2.24). Since $\langle{\mathfrak{A}}\rangle$ does not contain any of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$, none of these relations are primitively positively definable in $\operatorname{Feas}({\mathfrak{A}})$. By Theorem 5.8, $\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))\subseteq\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}^{\prime}))$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals and $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}^{\prime}))$ is in P. By Remark 2.9, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P. By Remark 2.24, $\operatorname{fPol}({\mathfrak{A}})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. Therefore, (4) holds. ∎
@@
Recall from Section 2.3 that a valued structure ${\mathfrak{A}}$ with $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$ has a (quantifier-free) first-order definition in $\operatorname{Aut}(\mathbb{Q};<)$ with the defining formulas being disjunctions of conjunctions of atomic formulas over $(\mathbb{Q};<)$. We continue by proving that the complexity dichotomy we gave in Theorem 5.23 is decidable, using the representation of ${\mathfrak{A}}$ by a first-order definition in $(\mathbb{Q};<)$ of the form described above.
Remark 5.24.
@@
We also obtain decidability if arbitrary first-order formulas may be used for defining the valued relations, because every first-order formula can be effectively transformed into such a formula. This holds more generally over so-called finitely bounded homogeneous structures; see, e.g., [36, Proposition 7]. Without the finite boundedness assumption, the problem can become undecidable [13].
Proposition 5.25.
@@
Given a first-order definition of a valued structure ${\mathfrak{A}}$ with a finite signature in $(\mathbb{Q};<)$, it is decidable whether $\operatorname{VCSP}({\mathfrak{A}})$ is in P or NP-complete.
Proof.
[…]
@@
We reformulate Theorem 5.23 with two mutually exclusive cases that capture the respective complexities of the VCSPs.
Corollary 5.26.
@@
  1. (1)
    $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$. In this case, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
  2. (2)
@@
Proof.
Let ${\mathfrak{A}}^{\prime}:=(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Theorem 5.8 states that either ${\mathfrak{A}}^{\prime}$ primitively positively defines one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, $\operatorname{Dis}$, or $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\operatorname{const}$, $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. Clearly, ${\mathfrak{A}}^{\prime}$ primitively positively defines a relation $R$ is and only if $R\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$, which is the case if and only if $R\in\langle{\mathfrak{A}}\rangle$.
It remains to discuss the implications for the complexity of $\operatorname{VCSP}({\mathfrak{A}})$. If (1) holds, then $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete by Theorem 5.23. On the other hand, if (1) does not hold, one of the cases (2)–(4) in Theorem 5.23 applies and $\operatorname{VCSP}({\mathfrak{A}})$ is in P. ∎
Note that the corollary above implies that if $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$, then the complexity of $\operatorname{VCSP}({\mathfrak{A}})$ is up to polynomial-time reductions determined by the complexity of the crisp relations ${\mathfrak{A}}$ can express. Loosely speaking, the complexity of such a VCSP is determined solely by the CSPs that can be encoded in this VCSP. We formulate an alternative and more concise variant of the previous result.
Corollary 5.27.
@@
  1. (1)
    ${\mathfrak{A}}$ pp-constructs $K_{3}$. In this case, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
  2. (2)

### v2

@@ 5.3 Expressibility of hard valued relations
It follows that $(<)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. Observe that $(<)_{0}^{1}$ equals $R_{1,0,1}$. Therefore, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ by Lemma 48, as we wanted to prove. If $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\min^{*}$, $\operatorname{mx}^{*}$, or $\operatorname{mi}^{*}$, we use the dual versions of Proposition 41 and Theorem 42 to analogously prove that $(>)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. Since $(<)_{0}^{1}(x,y)=(>)_{0}^{1}(y,x)$ we obtain $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ by Lemma 48. $\qed$
The last auxiliary lemma needed for the classification of temporal VCSPs shows that, assuming $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$, if an injective operation $f$ preserves all crisp relations expressible in $\mathfrak{A}$}, then $f$ is a fractional polymorphism of $\mathfrak{A}$} or $\operatorname{VCSP}({\mathfrak{A}})$ is NP-hard.
Lemma 53
Let $\mathfrak{A}$} be a valued structure with $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$ and $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$. Let $f\colon\mathbb{Q}^{2}\to\mathbb{Q}$ be an injective operation. Suppose that $f\notin\operatorname{fPol}({\mathfrak{A}})$ and $f\in\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Then $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$.
Proof 16
Let $R$ be a valued relation of $\mathfrak{A}$} of arity $k$ that is not improved by $f$. Then there exist $s,t\in\mathbb{Q}^{k}$ such that
$R(s)+R(t)<2R(f(s,t)).$
In particular, $R(s),R(t)<\infty$. Since $\operatorname{Feas}(R)\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ is improved by $f$, we have $R(f(s,t))<\infty$. Let $u:=f(s,t)$. Note that we must have $R(s)<R(u)$ or $R(t)<R(u)$. Moreover, $E_{u}=E_{s}\cap E_{t}$. Let $v\in\{s,t\}$ be such that $R(v)<R(u)<\infty$. Note that we have $E_{u}\subseteq E_{v}$. Let $O$ be a maximal subset of $O_{u}$ such that there exists $w\in\mathbb{Q}^{k}$ satisfying
  • 1.
@@
Since $R(w)\neq R(u)$ and $E_{u}\subseteq E_{w}$, there exist $i,j\in[k]$ such that $w_{i}\leq w_{j}$ and $u_{i}>u_{j}$. Without loss of generality me may assume $(i,j)=(1,2)$, because otherwise we permute the entries of $R$. Let $\phi$ be the following expression with the free variables $x_{1},x_{2}$.
$\min_{x_{3},\dots,x_{k}}\left(R(x_{1},\dots,x_{k})+\sum_{(p,q)\in E_{u}}(=)_{0}^{\infty}(x_{p},x_{q})+\sum_{(p,q)\in O}(<)_{0}^{\infty}(x_{p},x_{q})\right)$ (6)
Let $a,b\in\mathbb{Q}$ such that $a<b$. Then $\phi(b,a)=\phi(u_{1},u_{2})\leq R(u)$, because $O\subseteq O_{u}$. Suppose that $\phi(b,a)\leq R(w)$. Then there exists $w^{\prime}\in\mathbb{Q}^{k}$ such that $w_{1}^{\prime}>w^{\prime}_{2}$ and $w_{3}^{\prime},\dots,w_{k}^{\prime}$ realize the minimum for $\phi(b,a)$ in (6) and hence $\phi(b,a)=R(w^{\prime})\leq R(w)\leq R(v)$. In particular, the sums in (6) are finite. Therefore, $O\cup\{(2,1)\}\subseteq O_{w^{\prime}}$ and $E_{u}\subseteq E_{w^{\prime}}$. Since $(2,1)\in O_{u}\setminus O$, this contradicts the choice of $O$ and $w$. Therefore, $\phi(b,a)>R(w)$. Note that $\phi(w_{1},w_{2})\leq R(w)$. If $w_{1}<w_{2}$, then $\phi(a,b)\leq R(w)$ and $\phi$ expresses $R_{\alpha,\beta,\gamma}$ where $\beta=\phi(a,b)$ and $\gamma=\phi(b,a)$. In particular, $\beta<\gamma<\infty$. Therefore, by Lemma 49, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$. Otherwise we have $w_{1}=w_{2}$. Then $\phi(a,a)\leq R(w)$ and $\phi$ expresses $R_{\alpha,\beta,\gamma}$ where $\alpha=\phi(a,a)\leq R(w)<\phi(b,a)=\gamma$. If $\beta\geq\gamma$, then $\alpha<\min(\beta,\gamma)$, and otherwise $\beta<\gamma<\infty$. In both cases, $\operatorname{Cycl}\in\langle{\mathfrak{A}}\rangle$ by Lemma 49. $\qed$
5.4 Classification
We can now state and prove the complexity dichotomy for temporal VCSPs. We first phrase the classification with 4 cases, where we distinguish between the tractable cases that are based on different algorithms. As a next step, we formulate two corollaries each of which provides two concise mutually disjoint conditions that correspond to NP-completeness and polynomial-time tractability, respectively.
Theorem 54
@@ 5.4 Classification
    $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$ (see Definition 34), $T_{3}$, or $\operatorname{Dis}$ (see (4)). In this case, $\mathfrak{A}$} has a reduct $\mathfrak{A}$$^{\prime}$ over a finite signature such that $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
  2. 2.
    $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$.
  3. 3.
    $\operatorname{ll}\in\operatorname{fPol}({\mathfrak{A}})$ or $\operatorname{ll}^{*}\in\operatorname{fPol}({\mathfrak{A}})$.
  4. 4.
    $\pi_{1}^{2}\in\operatorname{fPol}({\mathfrak{A}})$ and $\operatorname{fPol}({\mathfrak{A}})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, or one of their duals.
@@
In cases 2–4, for every reduct $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$} over a finite signature, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P. Case 1 is mutually exclusive with each of the cases 2–4.
Proof 17
@@
Note that for every reduct $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$}, the automorphism group $\operatorname{Aut}({\mathfrak{A}}^{\prime})$ contains $\operatorname{Aut}({\mathfrak{A}})$ and hence is oligomorphic. If $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$, then there is a reduct $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$} over a finite signature such that $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is NP-hard by Lemma 4 and Theorem 44. By Theorem 21, $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in NP, therefore it is NP-complete. If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}}^{\prime})$ for every reduct $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$} over a finite signature, and $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P by Lemma 24. Suppose therefore that $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$ and that $\langle{\mathfrak{A}}\rangle$ does not contain any of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$.
@@
Let $\mathfrak{A}$$^{\prime}$ be a reduct of $\mathfrak{A}$} with a finite signature. If $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Sym}(\mathbb{Q})$, then by Theorem 32, $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})\subseteq\operatorname{fPol}({\mathfrak{A}}^{\prime})$ and $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P. Note that for $k\in\mathbb{N}$ and any $s,t\in A^{k}$, the tuple $\operatorname{inj}(s,t)$ lies in the same orbit as $\operatorname{ll}(s,t)$ under the action of $\operatorname{Aut}({\mathfrak{A}})$, hence $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$ implies $\operatorname{ll}\in\operatorname{fPol}({\mathfrak{A}})$ and therefore item 3 is satisfied.
@@
Suppose now that $\operatorname{Aut}({\mathfrak{A}})\neq\operatorname{Sym}(\mathbb{Q})$. By Lemma 50 we have $(<)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$, and hence $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$. By Lemma 52 we have that $\mathfrak{A}$} is essentially crisp, $\operatorname{ll}\in\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ or $\operatorname{ll}^{*}\in\operatorname{Pol}(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. If $\mathfrak{A}$} is not essentially crisp, by Lemma 53 we have that $\operatorname{ll}$ or $\operatorname{ll}^{*}$ lies in $\operatorname{fPol}({\mathfrak{A}})\subseteq\operatorname{fPol}({\mathfrak{A}}^{\prime})$. Then $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P by Lemma 47 and item 3 holds.
Finally, suppose that $\mathfrak{A}$} is essentially crisp. Then by Lemma 26 we have $\pi_{1}^{2}\in\operatorname{fPol}({\mathfrak{A}})$. Since $\operatorname{const}\not\in\operatorname{fPol}({\mathfrak{A}})$, we have $\operatorname{const}\not\in\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$ (see Remark 20). Since $\langle{\mathfrak{A}}\rangle$ does not contain any of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$, none of these relations are primitively positively definable in $\operatorname{Feas}({\mathfrak{A}})$. By Theorem 44, $\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))\subseteq\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}^{\prime}))$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals and $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}^{\prime}))$ is in P. Thus $\operatorname{VCSP}({\mathfrak{A}}^{\prime})$ is in P. By Remark 20, $\operatorname{fPol}({\mathfrak{A}})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. Therefore, item 4 holds.
To see that case 1 is mutually exclusive with cases 2–4, observe that none of the relations listed in item 1 is improved by any of the fractional polymorphisms $\operatorname{const}$, $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals, which proves the claim by Lemma 18. $\qed$
@@
Recall from Section 2.3 that a valued structure $\mathfrak{A}$} with $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$ has a (quantifier-free) first-order definition in $\operatorname{Aut}(\mathbb{Q};<)$ with the defining formulas being disjunctions of conjunctions of atomic formulas over $(\mathbb{Q};<)$. Using this representation of $\mathfrak{A}$}, we obtain the decidability of the complexity dichotomy from Theorem 54.
Remark 55
@@
We also obtain decidability if arbitrary first-order formulas may be used for defining the valued relations, because every first-order formula can be effectively transformed into such a formula. This holds more generally over so-called finitely bounded homogeneous structures; see, e.g., [39, Proposition 7]. Without the finite boundedness assumption, the problem can become undecidable [12].
Proposition 56
@@
Given a first-order definition of a valued structure $\mathfrak{A}$} with a finite signature in $(\mathbb{Q};<)$, it is decidable whether item 1 in Theorem 54 holds, in particular, whether $\operatorname{VCSP}({\mathfrak{A}})$ is in P or NP-complete.
Proof 18
Since by Theorem 54, item 1 in Theorem 54 is mutually exclusive with any of items 2–4, it is enough to show that it is decidable whether items 2–4 hold. Since $\mathfrak{A}$} has a finite signature, we can decide whether $\operatorname{const}$ improves $\mathfrak{A}$}, i.e., whether item 2 holds, by checking for each valued relation of $\mathfrak{A}$} whether the value of a constant tuple is minimal. Similarly, we can decide whether operations $\operatorname{ll}$ or $\operatorname{ll}^{*}$ improve $\mathfrak{A}$}: given a valued relation $R$ of $\mathfrak{A}$} of arity $k$, we need to choose finitely many tuples from each orbit of $A^{k}$ under the action of $\operatorname{Aut}({\mathfrak{A}})$ representing all distributions of values between non-positive and positive entries and test whether the relation is improved on pairs of these orbit representatives. Therefore, we can decide whether item 3 holds. Finally, we can decide whether $\pi_{1}^{2}$ improves $\mathfrak{A}$}, equivalently, whether $\mathfrak{A}$} is essentially crisp by Lemma 26. In this case $\operatorname{fPol}({\mathfrak{A}})$ contains $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals if and only if $\operatorname{Pol}(\operatorname{Feas}({\mathfrak{A}}))$ does (remark 20), which can be decided by Theorem 44. It follows that we can decide whether any of items 2–4 holds, which implies the statement. $\qed$
@@
We reformulate Theorem 54 with two mutually exclusive cases that capture the respective complexities of the VCSPs and relate it to the complexity of reducts of $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$.
Corollary 57
@@
  1. 1.
    $\langle{\mathfrak{A}}\rangle$ contains one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$. In this case, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
  2. 2.
@@
Proof 19
Let ${\mathfrak{A}}^{\prime}:=(\mathbb{Q};\langle{\mathfrak{A}}\rangle_{0}^{\infty})$. Theorem 44 states that either $\mathfrak{A}$$^{\prime}$ primitively positively defines one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, $\operatorname{Dis}$, or $\operatorname{Pol}({\mathfrak{A}}^{\prime})$ contains $\operatorname{const}$, $\min$, $\operatorname{mx}$, $\operatorname{mi}$, $\operatorname{ll}$, or one of their duals. Clearly, $\mathfrak{A}$$^{\prime}$ primitively positively defines a relation $R$ if and only if $R\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$, which is the case if and only if $R\in\langle{\mathfrak{A}}\rangle$.
It remains to discuss the implications for the complexity of $\operatorname{VCSP}({\mathfrak{A}})$. If item 1 holds, then $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete by Theorem 54. On the other hand, if item 1 does not hold, one of the items 2–4 in Theorem 54 applies and $\operatorname{VCSP}({\mathfrak{A}})$ is in P. $\qed$
Note that the corollary above implies that if $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$, then the complexity of $\operatorname{VCSP}({\mathfrak{A}})$ is up to polynomial-time reductions determined by the complexity of the crisp relations $\mathfrak{A}$} can pp-express. Loosely speaking, the complexity of such a VCSP is determined solely by the CSPs that can be encoded in this VCSP. We formulate an alternative and more concise variant of the previous result in the style of Theorem 45.
Corollary 58
@@
  1. 1.
    $\mathfrak{A}$}, or, equivalently, $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ $p$-constructs $K_{3}$. In this case, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
  2. 2.